ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 0 Kortesi, Peter, University of Miskolc, Hungary Junior Mathematical Congress - Challenge for future Mathematicians The History of the Junior Mathematical Congress has started in 1992, when in La Villete, Paris the first edition was held. The second edition was organized by Miskolc University in 1996 and the number of more then 340 participants has shown the increasing interest of young people and their teachers for similar meetings. The Executive Committee of the European Mathematical Society evaluated the Junior Mathematical Congress in Miskolc as a big success. It was decided that the next Junior Mathematical Congress should take place already in 2 years, in the summer of 1998. This time it is organized as a sattelite meeting of ICM 98 at Potsdam University, and beside its organizers there it is a pleasure to see those young people and colleagues, who where participants of the ’96 edition.